OpenStudy (anonymous):

Find the roots of the polynomial equation. x2 + 24 = 11x A. 12, 2 B. 3, 8 C. –8, –3 D. –12, –2

OpenStudy (anonymous):

Set equal to zero/ x^2 - 11x + 24 = 0 (x-8)(x-3) = 0 x = 8 and x = 3 (after setting each factor equal to 0). Choice (B) is correct.